# Prospect of GRB-Neutrino Detection with Enhanced Neutrino Detectors

### v1

@@ 3 The detection prospects of GRB 221009A-like events
It should be noted that, according to the conclusions of Ai & Gao (2023), for GRB 221009A, only if its internal shock originates from a region with a very large dissipation radius (a large variability timescale or a very large bulk Lorentz factor) can it match our expectation of not detecting neutrinos. Rudolph et al. (2023) presented a more sophysticated internal shock model, where the variability timescale is around $1$ s, and the final energy dissipation occurs at a radius of approximately $10^{16}\sim 10^{17}$ cm, which is even larger than that of the ICMART model. In this discussion, we adopt a conservative approach and assume that the minimum variability timescale is the classical $0.01$ s (Baerwald et al., 2011; Hümmer et al., 2012; Aartsen et al., 2017b). As shown in Figure 3, in the context of dissipative photosphere model, placing GRB 221009A at a redshift of 0.37 would still allow its generated neutrinos to be detectable by the current IceCube detector. With a less than twofold increase in the detector’s sensitivity, the neutrinos produced by GRB 221009A would remain observable even at a redshift of 0.5. In the context of the internal shock model, neutrinos from GRB 221009A can be detected within a redshift of $0.19$ without any increase in the detector’s effective area. A fivefold increase in the detector’s effective area would enable detection of neutrinos, provided the event occurs within a redshift of 0.5. In the case of the ICMART model, if the event occurs at the same redshift as GRB 221009A (z = 0.15), a tenfold increase in the detector’s effective area would be required to have a chance of detecting the neutrinos. It is worth noting that the above calculation are based on the effective area corresponding to the true declination angle of GRB 221009A. If the future event happens to occur at a declination where the detector’s effective area is maximized, the detection rate would increase significantly. In that case, only about 3 times increase of the detector’s sensitivity would be needed to detect the neutrinos from GRB 221009A-like bursts, at $z=0.15$ for ICMART model or at $z=0.5$ for internal shock model. The above conclusions are based on a 90% detection probability. Please note that the above discussion is based on the parameters used in Figure 2. The flux of neutrinos is influenced by these parameters, which may lead to potential uncertainties.
@@ 4 STACKED NEUTRINOS FROM LONG GRBs
Here, we use data from GRBweb [^2] For detailed data, see https://user-web.icecube.wisc.edu/~grbweb_public., an online platform that collects data from different telescopes, such as GBM (Hurley et al., 2013; von Kienlin et al., 2020), LAT (Ajello et al., 2019), Swift (Lien et al., 2016) and others. From 2019 to 2023, a total of 1503 gamma-ray bursts were recorded. We select those with fluence records and $T_{90}>2$ s for our calculations. If a source did not have a redshift measurement, we assigned a redshift value of $2.15$(Aartsen et al., 2017b). We adopt $E_{\text{break}}=200$ keV for all GRBs. For the bulk Lorentz factor ($\Gamma$) of GRBs, we derive it using the empirical relationship between the bulk Lorentz factor and the luminosity of the GRB (Liang et al., 2010; Lü et al., 2012; Zhang & Kumar, 2013), which is given by
@@
For simplicity, the uncertainties connected with this relation, which may result from the orientation of the jets, are not considered. We exclude GRB 210518C and GRB 230614C, even though their fluences were well recorded. This is because, in the absence of redshift measurements, assuming a redshift of 2.15 for these sources would result in an unreasonably high luminosity. As a result, they would produce a number of neutrinos comparable to those of the remaining $\sim 1,000$ gamma-ray bursts. In addition, GRB 221009A is also excluded from the samples.
@@
Figure 4: The solid lines are the predicted neutrino flux for the internal-shock, dissipative photosphere, and ICMART models. The dashed line represent the upper limit of observing the neutrino with a 90% probability for each model according to the effective areas of IceCube IC86-II. $\epsilon_{p}/\epsilon_{e}=3$, $\epsilon_{B}/\epsilon_{e}=1$, $\delta t_{\rm min}=0.01$ s, $R_{\rm ICMART}=10^{15}$ cm are adopted.
@@
We assume that GRBs will continue to be observed in the coming years at the same detection rate as during the period from 2019 to 2023. Using the same parameters as those applied in the models shown in Figure 4, we calculate the evolution of the probability of detecting neutrinos from GRBs over the detector’s operational time. The results are shown in Figure 5. Assuming the detector’s effective area remains unchanged, the dissipative photosphere model and the internal shock model require 4.35 years and 7.11 years, respectively, to achieve a $90\%$ detection probability. In contrast, the ICMART model can only reach a detection probability of $58\%$, even with an accumulation time of 100 years.
@@
Figure 5: For different models, the detection probability of neutrinos varies with the accumulation time. The solid line represents the current effective area of IceCube, while the dashed and dotted lines represent the effective area expanded by a factor of 5 and 10, respectively. The dotted lines corresponding to the dissipative photosphere and internal shock models have been bolded for better visibility. $\epsilon_{p}/\epsilon_{e}=3$, $\epsilon_{B}/\epsilon_{e}=1$, $\delta t_{\rm min}=0.01$ s, $R_{\rm ICMART}=10^{15}~{}{\rm cm}$ are adopted.
@@
With next-generation neutrino detectors, the increase in the effective area will significantly enhance the probability of jointly detecting GRBs and neutrinos. If the detector’s effective area is increased fivefold relative to IceCube IC86-II, the dissipative photosphere and internal shock models would require only one year of accumulation to achieve detection probabilities of 80% and 93%, respectively. In contrast, the ICMART model would require 52 years of accumulation to reach a 90% detection probability. With a tenfold expansion of the detector’s effective area, the dissipative photosphere and internal shock models would quickly approach a 100% detection probability. However, even with 10 years of accumulation,the ICMART model would only achieve a 58% detection probability.
@@
On the other hand, even if the neutrino counterparts of GRBs remain undetected, much more stringent constraints can be placed on the free parameters of different GRB models. If the parameters are constrained to an unacceptable range, it can be concluded that the corresponding GRB model can be rule out. Inspired by observations of GRBs and their afterglows, we assume reasonable parameters to be $\epsilon_{p}/\epsilon_{e}>1$, $\epsilon_{B}/\epsilon_{e}<1$ (Gao et al., 2015). And we still adopt $\delta t_{\rm min}=0.01$ s and $R_{\rm ICMART}=10^{15}$ cm. In the future, assuming the events accumulated over five years, similar to those from 2019 to 2023, using the non-detection results from an enhanced neutrino detector and the parameter space shown in Figure 6, one may conclude that:
@@
  • For the internal shock model, if the effective area has been increased by a factor of $5.5$ relative to IceCube IC86-II, then for $\delta t_{\rm min}=0.01$ s, $\epsilon_{p}/\epsilon_{e}$ must be less than approximately $0.96$. This implies that the internal shock model with $\delta t_{min}=0.01$s is not generally applicable to GRBs. In addition to the conservative case, some also suggest that the minimum variability timescale could be 0.1 s (Zhang & Kumar, 2013). We have also considered this scenario and find that if the detector’s effective area is increased by a factor of 19 and neutrinos are still not detected, then for $\delta t_{\rm min}=0.1$ s, $\epsilon_{p}/\epsilon_{e}$ must be less than 1. Therefore, to rule out this scenario, the detector’s effective area would need to be increased by a factor of 19.
@@ 5 Conclusions and Discussions
Thus, the lack of neutrinos associated with GRB 221009A is consistent with implications from multiband EM observations suggesting that a magnetically dominated jet was launched. With future enhanced neutrino detectors, if the effective area is approximately $10$ times larger than that of IceCube IC86-II, we would be able to detect neutrinos from such a GRB event which have the same redshit with GRB 221009A, even if produced under the ICMART model. If we are particularly lucky, and the event occurs at a declination corresponding to the effective area of maximum detector efficiency, then increasing the effective area by a factor of $3$ would be sufficient to detect the neutrinos it produces.
We then calculated the cumulative neutrino flux from stacked GRBs and analyzed 1,142 sources from 2019 to 2023. We considered a scenario where future detectors with an increased effective area observing these $1,142$ sources over a 5-year period. If the effective area is increased $4$ times relative to IceCube IC86-II and no neutrinos are detected, the general applicability of the dissipative photosphere model would be strongly questioned. If expanded $5.5$ times, the same issue appears to the internal shock model. For the ICMART model, even if the detector’s effective area is increased by a factor of $10$ and no associated neutrinos are detected, the model can still survive.
Here are three cautions: (1) For the internal shock model, we assume the minimum variability timescale of the GRB light curve is $\delta t_{\rm min}\sim 0.01$ s, which is a classical theoretical value. However, if $\delta t_{\rm min}$ is much greater (like $0.1$ s inferred from some of observed minimum variability timescale (Golkhou et al., 2015; Camisasca et al., 2023)), the internal shock model should have a radiation radius comparable to that of the ICMART model (Rudolph et al., 2023), making neutrinos production in the internal shock model also very inefficient. (2) When we rule out models using stacked GRB observations, we mean ruling out the possibility that a single model applies to all GRBs. In fact, there may be multiple channels responsible for producing GRBs. (3) Our discussion is valid only in the “one-zone” scenario, where protons are accelerated in the same region where the gamma-ray photons are emitted.
Studies predict that low-luminosity GRBs might be more efficient generators of high-energy neutrinos (Murase et al., 2006b; Gupta & Zhang, 2007). Similarly, short GRBs with relatively lower bulk Lorentz factors in their jets could also be potential sources of high-energy neutrinos (Rudolph et al., 2024). Currently operating powerful gamma-ray and X-ray detectors could detect more of these relatively faint events, thereby providing better constraints on GRB models.
@@
Appendix A Appendix A
The cross section for $p\gamma$ interaction of photons in the proton’s rest frame are taken from Yu et al. (2008) and shown in Figure 7.

### v2

@@ 3 The detection prospects of GRB 221009A-like events
It should be noted that, according to the conclusions of Ai & Gao (2023), for GRB 221009A, only if its internal shock originates from a region with a very large dissipation radius (a large variability timescale or a very large bulk Lorentz factor) can it match our expectation of not detecting neutrinos. Rudolph et al. (2023) presented a more sophysticated internal shock model, where the variability timescale is around $1$ s, and the final energy dissipation occurs at a radius of approximately $10^{16}\sim 10^{17}$ cm, which is even larger than that of the ICMART model. In this discussion, we adopt a conservative approach and assume that the minimum variability timescale is the classical $0.01$ s (Baerwald et al., 2011; Hümmer et al., 2012; Aartsen et al., 2017b). As shown in Figure 3, in the context of the dissipative photosphere model, placing GRB 221009A at a redshift of 0.37 would still allow its generated neutrinos to be detectable by the current IceCube detector. With a less than twofold increase in the detector’s sensitivity, the neutrinos produced by GRB 221009A would remain observable even at a redshift of 0.5. In the context of the internal shock model, neutrinos from GRB 221009A can be detected within a redshift of $0.19$ without any increase in the detector’s effective area. A fivefold increase in the detector’s effective area would enable detection of neutrinos, provided the event occurs within a redshift of 0.5. In the case of the ICMART model, if the event occurs at the same redshift as GRB 221009A (z = 0.15), a tenfold increase in the detector’s effective area would be required to have a chance of detecting the neutrinos. It is worth noting that the above calculations are based on the effective area corresponding to the true decl angle of GRB 221009A. If a future event occurs at a decl where the detector’s effective area is maximized (i.e., near the equator), the detection rate would increase significantly. This maximized effective area is obtained by integrating over the predicted neutrino energy range from $10^{2}$ GeV to $10^{9}$ GeV. In that case, only about $3$ times the increase of the detector’s sensitivity would be needed to detect the neutrinos from GRB 221009A-like bursts, at $z=0.15$ for the ICMART model or at $z=0.5$ for the internal shock model. The above conclusions are based on a 90% detection probability. Please note that the above discussion is based on the parameters used in Figure 2. The flux of neutrinos is influenced by these parameters, which may lead to potential uncertainties.
@@ 4 STACKED NEUTRINOS FROM LONG GRBs
Here, we use data from GRBweb [^1] For detailed data, see https://user-web.icecube.wisc.edu/~grbweb_public., an IceCube project that collects data from different telescopes, such as GBM (Hurley et al., 2013; von Kienlin et al., 2020), LAT (Ajello et al., 2019), Swift (Lien et al., 2016) and others. From 2019 to 2023, a total of 1503 gamma-ray bursts were recorded. We select those with fluence records and $T_{90}>2$ s for our calculations. If a source did not have a redshift measurement, we assigned a redshift value of $2.15$(Aartsen et al., 2017b). We adopt $E_{\text{break}}=200$ kev, $\alpha=1$ and $\beta=2$ for all GRBs. For the bulk Lorentz factor ($\Gamma$) of GRBs, we derive it using the empirical relationship between the bulk Lorentz factor and the luminosity of the GRB (Liang et al., 2010; Lü et al., 2012; Zhang & Kumar, 2013), which is given by
@@
For simplicity, the uncertainties connected with this relation, which may result from the orientation of the jets, are not considered. We exclude GRB 210518A and GRB 230614C, even though their fluences were well recorded in GRBweb. This is because, in the absence of redshift measurements, assuming a redshift of 2.15 for these sources would result in an unreasonably high luminosity. As a result, they would produce a number of neutrinos comparable to those of the remaining $\sim 1000$ gamma-ray bursts. In addition, GRB 221009A is also excluded from the samples.
@@
Figure 4: The solid lines are the predicted neutrino flux for the internal shock, dissipative photosphere, and ICMART models. The dashed lines represent the upper limit of observing the neutrino with a 90% probability for each model according to the effective areas of IceCube IC86-II. $\epsilon_{p}/\epsilon_{e}=3$, $\epsilon_{B}/\epsilon_{e}=1$, $\delta t_{\rm min}=0.01$ s, and $R_{\rm ICMART}=10^{15}$ cm are adopted.
@@
We assume that GRBs will continue to be observed in the coming years at the same detection rate as during the period from 2019 to 2023. Using the same parameters as those applied in the models shown in Figure 4, we calculate the evolution of the probability of detecting neutrinos from GRBs over the detector’s operational time. The results are shown in Figure 5. Assuming the detector’s effective area remains unchanged, the dissipative photosphere model and the internal shock model require 4.35 yr and 7.11 yr, respectively, to achieve a $90\%$ detection probability. In contrast, the ICMART model can only reach a detection probability of $58\%$, even with an accumulation time of 100 yr.
@@
Figure 5: For different models, the detection probability of neutrinos varies with the accumulation time. The solid line represents the current effective area of IceCube, while the dashed and dotted lines represent the effective area expanded by a factor of 5 and 10, respectively. The dotted lines corresponds to the dissipative photosphere, and the internal shock models have been bolded for better visibility. $\epsilon_{p}/\epsilon_{e}=3$, $\epsilon_{B}/\epsilon_{e}=1$, $\delta t_{\rm min}=0.01$ s, and $R_{\rm ICMART}=10^{15}~{}{\rm cm}$ are adopted.
@@
With next-generation neutrino detectors, the increase in the effective area will significantly enhance the probability of jointly detecting GRBs and neutrinos. If the detector’s effective area is increased fivefold relative to IceCube IC86-II, the dissipative photosphere and internal shock models would require only 1 yr of accumulation to achieve detection probabilities of 80% and 93%, respectively. In contrast, the ICMART model would require 52 yr of accumulation to reach a 90% detection probability. With a tenfold expansion of the detector’s effective area, the dissipative photosphere and internal shock models would quickly approach a 100% detection probability. However, even with 10 years of accumulation, the ICMART model would only achieve a 58% detection probability.
@@
On the other hand, even if the neutrino counterparts of GRBs remain undetected, much more stringent constraints can be placed on the free parameters of different GRB models. If the parameters are constrained to an unacceptable range, it can be concluded that the corresponding GRB model can be ruled out. Inspired by observations of GRBs and their afterglows, we assume reasonable parameters to be $\epsilon_{p}/\epsilon_{e}>1$, $\epsilon_{B}/\epsilon_{e}<1$ (Gao et al., 2015). And we still adopt $\delta t_{\rm min}=0.01$ s and $R_{\rm ICMART}=10^{15}$ cm. In the future, assuming the events accumulated over 5 yr, similar to those from 2019 to 2023, using the nondetection results from an enhanced neutrino detector and the parameter space shown in Figure 6, one may conclude that:
@@
  • For the internal shock model, if the effective area has been increased by a factor of $5.5$ relative to IceCube IC86-II, then for $\delta t_{\rm min}=0.01$ s, $\epsilon_{p}/\epsilon_{e}$ must be less than approximately $0.96$. This implies that the internal shock model with $\delta t_{min}=0.01$s is not generally applicable to GRBs. In addition to the conservative case, some also suggest that the minimum variability timescale could be 0.1 s (Zhang & Kumar, 2013). We have also considered this scenario and find that if the detector’s effective area is increased by a factor of 19, and neutrinos are still not detected, then for $\delta t_{\rm min}=0.1$ s, $\epsilon_{p}/\epsilon_{e}$ must be less than 1. Therefore, to rule out this scenario, the detector’s effective area would need to be increased by a factor of 19.
@@
5 Predictions with upcoming neutrino detectors
A number of innovative neutrino detectors have been proposed, and several are presently being constructed. With their proposed sensitivity curves, one can estimate the magnification factors with respect to IceCube IC86-II. Here we assume the neutrino spectrum follows $E^{-2}$. By integrating over the energy range from $10^{2}$ GeV to $10^{9}$ GeV using this spectrum, we obtain IceCube IC86-II’s sensitivity at different declinations. Since IceCube’s detection capability is highly sensitive to decl, we take the average sensitivity across different values of the sine of the decl and compare this average with the sensitivities of other detectors, which are shown in the corresponding references. We find that the magnification factors are $\sim 8$ for IceCube Gen2 (Omeliukh et al., 2022), $\sim 10$ for KM3NeT (Aiello et al., 2019), and $\sim 30$ for TRIDENT (Ye et al., 2024). Comparing these magnification factors with the requirements shown in Figure 3, we find that all future neutrino detectors can achieve an ideal detection prospect for a single source resembling GRB 221009A, provided the source occurs at the same redshift. IceCube Gen2, with a relatively low magnification factor, may not be fully located within the detectable parameter space for all GRB prompt emission models. For the stacked neutrinos from GRBs, based on the general results shown in Section 4, we explicitly outline the conditions required to support or effectively rule out different GRB models in Figure 7, if no neutrino-GRB association is confirmed, with the proposed neutrino detectors’ effective areas marked. All three proposed neutrino detectors have the capability to rule out the dissipative photosphere model and the internal shock model with $\delta t_{\rm min}=0.01s$, over a 5 yr period of operation and can very likely achieve this progress within 2 to 3 yr. For the internal shock model with $\delta t_{\rm min}=0.1s$, only TRIDENT has the capability to rule it out within 5 yr (specifically in 3.17 yr). For the ICMART model, the required magnification factor is far beyond the capabilities of both current and near-future detectors. However, Agostini et al. (2020) proposed that multiple neutrino detectors can be combined into a network, which would significantly enhance detection capabilities and potentially meet the requirement to rule out the ICMART model.
@@ 5 Predictions with upcoming neutrino detectors
Figure 7: The required magnification factors and data accumulation periods for future neutrino detectors to effectively rule out different GRB prompt emission models are shown. The upper-right region of each line represents the required parameter space. The three dots indicate the required magnification factors of the detectors’ effective areas, relative to IceCube IC86-II, to exclude the dissipative photosphere model and the internal shock model ($\delta t_{\rm min}=0.01s$ or $\delta t_{\rm min}=0.1s$) under a five-year observation. The three vertical black dashed lines represent the magnification factors of IceCube Gen2, KM3NeT, and TRIDENT.
@@ 6 Conclusions and Discussions
Thus, the lack of neutrinos associated with GRB 221009A is consistent with implications from multiband EM observations suggesting that a magnetically dominated jet was launched. With future enhanced neutrino detectors, if the effective area is approximately $10$ times larger than that of IceCube IC86-II, we would be able to detect neutrinos from such a GRB event that has the same redshift with GRB 221009A, even if produced under the ICMART model. If we are particularly lucky, and the event occurs at a decl corresponding to the effective area of maximum detector efficiency, then increasing the effective area by a factor of $3$ would be sufficient to detect the neutrinos it produces.
We then calculated the cumulative neutrino flux from stacked GRBs and analyzed 1142 sources from 2019 to 2023. We considered a scenario where future detectors with an increased effective area observe these 1142 sources over a 5 yr period. If the effective area is increased $4$ times relative to IceCube IC86-II and no neutrinos are detected, the general applicability of the dissipative photosphere model would be strongly questioned. If expanded $5.5$ times, the same issue appears in the internal shock model. For the ICMART model, even if the detector’s effective area is increased by a factor of $10$ and no associated neutrinos are detected, the model can still survive. We also compare the above results with the capabilities of upcoming neutrino detectors and find that they can probe the dissipative photosphere model and the internal shock model. However, a significant gap still remains for the ICMART model.
Please note the underlying conditions required to derive these results: (1) For the internal shock model, we assume the minimum variability timescale of the GRB light curve is $\delta t_{\rm min}\sim 0.01$ s, which is a classical theoretical value. However, if $\delta t_{\rm min}$ is much greater (like $0.1$ s inferred from some of observed minimum variability timescale (Golkhou et al., 2015; Camisasca et al., 2023)), the internal shock model should have a radiation radius comparable to that of the ICMART model (Rudolph et al., 2023), making neutrino production in the internal shock model also very inefficient. (2) When we rule out models using stacked GRB observations and predicted effective areas from future neutrino detectors, we mean ruling out the possibility that a single model applies to all GRBs. In fact, there may be multiple channels responsible for producing GRBs. (3) Our discussion is valid only in the “one-zone” scenario, where protons are accelerated in the same region where the gamma-ray photons are emitted.
Studies predict that low-luminosity GRBs might be more efficient generators of high-energy neutrinos (Murase et al., 2006; Gupta & Zhang, 2007). Similarly, short GRBs with relatively lower bulk Lorentz factors in their jets could also be potential sources of high-energy neutrinos (Rudolph et al., 2024). Currently operating powerful gamma-ray and X-ray detectors could detect more of these relatively faint events, thereby providing better constraints on GRB models.
acknowledgments
@@
Appendix A Formulae for calculating neutrino flux
Equation 2 shows the general formula for calculating the flux of neutrinos from a single GRB source. Here, we describe the details of each term in the equation. The maximum proton energy $E_{\rm p,max}$ could be estimated by equaling the dynamical timescale $t^{\prime}\approx R/\Gamma c$ with the accelerating timescale of protons $t^{\prime}_{\rm acc}\approx E^{\prime}_{p}/\left(eB^{\prime}c\right)$, where $\Gamma$ is bulk motion Lorentz factor of the GRB jet. The $\prime$” denotes physical quantities defined in the jet’s comoving frame, and hereafter. Hence, we have
$E_{p,\text{max}}\leq 7.59\times 10^{11}\text{ GeV}\left(\frac{\epsilon_{B}}{\epsilon_{e}}\right)^{1/2}L_{\text{GRB},52}^{1/2}\Gamma_{2.5}^{-1}$ (A1)
$E_{p,\text{min}}>\Gamma m_{p}c^{2}=2.56\times 10^{2}\text{ GeV }\Gamma_{2.5}.$ (A2)
[…]
where $\gamma_{p}$ is the random-motion Lorentz factor of proton in the comving frame and $\bar{\epsilon}_{\gamma}$ is the energy of photo in the proton’s rest frame. $\kappa_{p\gamma}\approx 0.2$ is the inelasticity coefficient of the photon in the proton’s rest frame. $\sigma_{p\gamma}(\bar{\epsilon}_{\gamma})$ is the cross section of the $p\gamma$ interaction in the proton’s rest frame. Here, we do not treat the cross section for the $p\gamma$ interaction as a constant but rather as a function of the photon energy $\epsilon_{\gamma}$, with the values in the proton’s rest frame taken from (Yu et al., 2008) and shown in Figure 8. $n_{\gamma}$ is the spectrum of GRBs in the jet’s comoving frame, which is assumed to be in the form of a band spectrum (Poolakkil et al., 2021). Then, we can calculate the generation efficiency of pions as (Waxman & Bahcall, 1997; Kimura, 2022)
$f_{p\gamma}\approx 1-\exp(-{t_{p\gamma}^{-1}}/{t_{\rm dyn}^{-1}})\approx{\rm min}({t_{p\gamma}^{-1}}/{t_{\rm dyn}^{-1}},1).$ (A4)
The synchrotron cooling of $\pi^{+}$($\mu^{+}$) would have a suppressive effect on the production of neutrinos. The decay timescale of $\pi^{+}$ is $t^{\prime}_{\pi^{+},\rm dec}=2.6\times 10^{-8}\rm{s}~{}\gamma_{\pi^{+}}$ (e.g. Kimura, 2022) , where $\gamma_{\pi^{+}}$ is the Lorentz factor for $\pi^{+}$ in the jet’s comoving frame. For the relativistic $\pi^{+}$, the synchrotron cooling timescale can be calculated as $t^{\prime}_{\pi,\text{syn}}={6\pi m_{\pi^{+}}c}/{\gamma_{\pi}\sigma_{T,\pi^{+}}B^{\prime 2}}$ where $m_{\pi^{+}}=0.15m_{p}$ is the rest mass of $\pi^{+}$. The Thomson scattering cross section of $\pi^{+}$ can be estimated from that of the electrons as $\sigma_{T,\pi^{+}}=\left(m_{e}/m_{\pi^{+}}\right)^{2}\sigma_{T,e}$. $\gamma_{\pi^{+}}$ corresponds to the energy of neutrinos to be generated. Since the energy of $\pi^{+}$ would be shared nearly equally by four leptons, we obtain $E_{\nu}=\frac{1}{4}D\gamma_{\pi}^{+}m_{\pi^{+}}c^{2}$, where $D\approx\Gamma$ is the Doppler factor.  [^2] We consider the case that the observational angle is $1/\Gamma$, where the emissivity of GRBs reaches the maximum.. The cooling factor is then written as
$f_{\rm cooling}\approx 1-\exp(-({t_{\rm syn}^{-1}+t_{\rm dyn}^{-1}})/{t_{\rm dec}^{-1}}).$ (A5)